OpenStudy (anonymous):

Find derivative using the power rule cube root of X - cube root of X^2

OpenStudy (anonymous):

\[\sqrt[3]{x}=x^{\frac{1}{3}}\] and \[\sqrt[3]{x^2}=x^{\frac{2}{3}}\] so now can use power rule

OpenStudy (anonymous):

\[f(x)=x^{\frac{1}{3}}-x^{\frac{2}{3}}\] \[f'(x)=\frac{1}{3}x^{\frac{1}{3}-1}-\frac{2}{3}x^{\frac{2}{3}-1}\]

OpenStudy (anonymous):

\[f'(x)=\frac{1}{3}x^{-\frac{2}{3}}-\frac{2}{3}x^{-\frac{1}{3}}\]

OpenStudy (anonymous):

and finally \[f'(x)=\frac{1}{3\sqrt[3]{x^2}}-\frac{2}{3\sqrt[3]{x}}\]

OpenStudy (anonymous):

back in radical form

OpenStudy (anonymous):

hope all the gory details were complete and clear

OpenStudy (anonymous):

and correct!

OpenStudy (anonymous):

I'm sure it was. I'd hate to see you go through all that trouble for nothing